A crew rowing with the current traveled 16 miles in 2 hours; against the current, the crew rowed 8 miles in 2 hours. Find the rate at which the crew is rowing in still water and the rate of the current.
Answer by josmiceli(19441)

s = crew's speed in still water
c = speed of current
s + c = 16/2
s - c = 8/2
add the equations
2s = 24/2
s = 6
crew's speed in still water = 6 mi/hr
s + c = 16/2
6 + c = 8
c = 8 - 6
c = 2
speed of current = 2 mi/hr